OpenStudy (anonymous):

Solve: 3|a+d|+b when b = 3

OpenStudy (anonymous):

I originally had: |-2+-6|+3 = |-2 + -6| + 3 = |-8| + 3 = 8+3 = 11 but my teacher said it was wrong because i forgot to multiply the three in front ><

hartnn (hartnn):

so, 3(8)+3 = 27

hartnn (hartnn):

whats was the doubt ? :O

OpenStudy (anonymous):

is that it....? My teacher said something about an 8 but it confused me.. the rest is fine, all i have to do is 3(8)+3? (which = 27 :P) @hartnn

hartnn (hartnn):

it seems thats it....but can you plz post the entire question exactly, so that i can be sure.

OpenStudy (anonymous):

I did.....@hartnn

hartnn (hartnn):

you just wrote "Solve: 3|a+d|+b when b = 3" there's no equation to solve....

OpenStudy (anonymous):

3|a+d|+b a = -2, b = 3, c = -4 and d = -6

hartnn (hartnn):

3 |-2+-6| + 3 = 3 |-8|+3 = 3*8+3 =27 then yes, thats it! :)

OpenStudy (anonymous):

Thanks, Goku. :>

hartnn (hartnn):

Welcome ^_^